OpenStudy (anonymous):

Number 14 on section 5.5 of the book. " calculus. Early transcendentals tenth edition"

OpenStudy (anonymous):

Can you copy the question to here?

OpenStudy (anonymous):

@RyanRyan

OpenStudy (abb0t):

Is it an integration problem?

OpenStudy (anonymous):

It's an integration problem. " sketch the region whose signed area is represented by the definite integral, and evaluate the integral using the approximate formula from geometry, where needed ".. Integral sign with a 2 on top and 0 on bottom ( 1-1/2x)dx

OpenStudy (anonymous):

Yes

OpenStudy (abb0t):

\[\int\limits_{0}^{2} (1-\frac{1}{2}x)dx = \frac{1}{2} \times 2 \times 1 = 1\]

OpenStudy (anonymous):

yes abb0t, that is it. I am having to do a,b,c, and d.

OpenStudy (abb0t):

trapezoid is the same thing, except evaluated from -1<x<1 with base 2 and parrallel sides 1.5 and .5 so it's \[\int\limits_{-1}^{1}(1-\frac{1}{2}x)dx\]

OpenStudy (abb0t):

triangle is \[\int\limits_{2}^{3}(1-\frac{1}{2})dx = 1 \times 0.5\] and last one is \[\int\limits_{0}^{3}(1-\frac{1}{2})dx \] which you can do yourself.

OpenStudy (anonymous):

thanks, how did you get the answers so quick? Are you using the Riemann Summ??

OpenStudy (abb0t):

No. \[\int dx - \int (\frac{1}{2}x)dx\] i separated them \[\int_{a}^{b} dx = x,~~evaluated~ ~from~~a<x<b\] and you know that by the fundamental theorem of calculus, that you simply use the difference. then for the second integral \[\int_{a}^{b} \frac{1}{2}xdx = \frac{1}{4}x^2\] also evaluated from a<x<b where a and b are your bounds.

OpenStudy (anonymous):

thanks